# PerfTracker: Online Performance Diagnosis for Large Model Training

### v3

@@ 4. PerfTracker Design
PerfTracker is an online performance troubleshooting system for large-scale model training (LMT). It embodies the insights discussed in §3 in a production service for our customers who run LMT on our GPU clusters. PerfTracker employs a highly efficient approach to detecting performance degradation without continuous online profiling, introducing little overhead on LMTs. When a performance degradation is detected, PerfTracker automatically troubleshoots the performance issue. PerfTracker uses online profiling to collect fine-grained information, but it does not try to aggregate raw profile data of every function; instead, PerfTracker summarizes the runtime behavior patterns of each function at every worker node and aggregates only the concise behavior patterns. The diagnosis is performed on the aggregated behavior patterns based on the differential observability principle.
@@
Figure 6 gives an overview of PerfTracker. PerfTracker has three main components: (1) detecting performance degradation of LMT to trigger online profiling, (2) summarizing runtime behavior patterns of each function from raw profiling data, and (3) a centralized localization algorithm that pinpoints the root-cause function based on the behavior patterns. Both (1) and (2) are deployed per worker, while (3) is global.
@@
Usage. To deploy PerfTracker, GPU cluster providers do not need access to the user code. The user only needs to import PerfTracker in the LMT code (import PerfTracker); no other code change is needed. Behind the line of import, PerfTracker initializes the detector of performance degradation, registers handler functions to start profiling, and creates a daemon process for summarizing and uploading behavior patterns. Figure 7 illustrates the output of PerfTracker. As a function-centric approach, PerfTracker pinpoints which functions on which workers are executed abnormally, and describes how they behave differently from expectation or from other workers. In Figure 7, all workers encounter slow socket receive in data loading, indicating slow storage I/O. Worker 7 is slow in collective communication, indicating a network degradation on its connected link. In addition, worker 0,1,2, and 3 execute GEMM (a CUDA kernel function (gem, 2024)) slowly because of not fully utilizing GPUs, indicating the problem of GPU throttling.
Figure 7. An example of PerfTracker output.
4.1. Detecting Performance Degradation
@@
Profiling an LMT task online introduces overhead. To minimize the overhead while collecting profiling data timely when performance issues occur, PerfTracker automatically detects performance degradation and then makes profiling.
@@
For an LMT task, the time spent on one training iteration is the key performance metric. PerfTracker devises a solution to monitor the training time of each iteration without accessing user code, and then triggers online profiling if the iteration time is abnormal. Note that approaches that require accessing user data (such as logs) do not work with our usage model.
Figure 8. PerfTracker’s performance degradation detection.
@@
Indicators of iteration time. A PyTorch training iteration always involves several dataloader.next() calls, followed by several optimizer.step() calls (the number depends on training parameters like pipeline parallelism). In this sequence of events, the duration from the first dataloader.next() to the last optimizer.step() is regarded as the duration of a complete training iteration.
@@
When it is import-ed, PerfTracker wraps the two functions by adding time counters. Basically, PerfTracker replaces the two PyTorch functions with the wrapped versions. Since the two functions are Python functions (which are not compiled before LMT executions), the replacement is done at runtime.
@@
Iteration detection. During an LMT, PerfTracker continuously records event sequences of dataloader.next() and optimizer.step(). After detecting $M$ (=10 in practice) identical sequences starting with dataloader.next() and ending with optimizer.step(), this sequence is defined as the training iteration sequence.
@@
Performance degradation detection. PerfTracker continuously monitors incoming events of dataloader.next() and optimizer.step() with the identified training iteration sequence. Each time PerfTracker successfully matches a complete training iteration sequence, it records the duration of that iteration. PerfTracker considers that performance degradation has occurred in the following two situations: (1) The average duration of the recent $N$ (=50 in practice) iterations exceeds the recent shortest iteration time by more than 5%. (2) The current training iteration sequence has not yet been fully matched, but the time elapsed since the last event received is at least 5×5$\times$5 × the average iteration duration (indicating the training is blocked). If PerfTracker fails to match a training iteration after $K$ (=200 in practice) consecutive event receptions, it goes back to the previous iteration detection phase to redetect the training iteration sequence. While this does not often happen in common LMT programs, it significantly enhances the robustness of PerfTracker when users implement special functionalities in their code, to make sure the algorithm always works. Figure 8 illustrates the procedure.
Profiling data generation. In PerfTracker, each LMT worker is connected to a PerfTracker daemon. Upon detection of performance degradation, all daemons are notified via TCP sockets. Each daemon sends a signal to the corresponding worker. The worker receives the signal and starts profiling via a handler function (registered when PerfTracker is imported). This ensures that profiling is executed in the main thread of LMT (or else some APIs may fail, like CUPTI in Torch Profiler).
@@
A profiling session lasts for 20 seconds by default (which is configurable). We apply Torch Profiler to obtain function execution events for Python functions, CPU operations, memory operations, and CUDA kernels. We utilize nsys to sample hardware metrics at 10 kHz, such as GPU, DRAM, NVLink, PCIe, and the network. The profiling data is then collected by the PerfTracker daemon of each worker.
4.2. Summarizing Behavior Patterns
@@
PerfTracker summarizes behavior patterns of each function at the worker level. Note that PerfTracker only concerns functions on the critical path of LMT (Figure 9)—only these functions contribute to the end-to-end performance, despite the huge number of function executions in an LMT.
Figure 9. The critical path of LMT performance.
@@
PerfTracker’s solution is (1) finding the function execution events on the critical path, including GPU computation kernels, collective communication functions, memory operations, Python functions, and all other functions executed in LMT; and (2) clustering all execution events of each function (for Python functions, the entire call stack must be identical to be considered the same function), then defining several patterns to summarize the behavior of each function.
@@
Critical path. As shown in Figure 9, different types of functions in LMT have different priorities: GPU computation kernel functions have the highest priority; memory operations (e.g., malloc, memcpy) have second priority; collective communication kernel functions (e.g., NCCL AllReduce) have third priority; and Python functions have fourth priority. A function’s execution (or a subinterval of it) is on the LMT’s critical path if and only if no higher-priority function is executing during that time. For Python functions, we additionally require that the function belongs to the training thread of the LMT (i.e., not a thread initiated by _bootstrap), and that no child functions called by this function are executing.
The rationale for such a definition is that, in a well-optimized LMT, the majority of time should be spent on GPU computations and their associated memory operations. Collective communications and Python functions should try to overlap with GPU computations as much as possible. In performance diagnosis, we only focus on the non-overlapping parts.
@@
(3) $C(f,w,t)=\begin{cases}1&\text{$f$ is on $w$'s critical path at time $t$},\\ 0&\text{Otherwise }.\end{cases}$
@@
$\mu_{f,w}$ represents the average hardware resource utilization during all executions of function $f$ on worker $w$. In this context, “resource” refers to hardware resources that determine performance of function $f$. For example, GPU computation kernels correspond to GPU SM utilization, Python functions to CPU utilization, intra-worker collective communication to NVLink utilization, and inter-worker collective communication corresponds to the bandwidth utilization between GPU and NIC, due to the significantly lower NIC bandwidth compared to intra-worker NVLink bandwidth.
(4) $\mu_{f,w}=\frac{\sum_{e\in E_{f,w}}{|L(e)|avg_{U}(L(e))}}{\sum_{e\in E_{f}}{|L(e)|}}$
@@
Figure 10. Critical duration of a collective communication function. A worker enters the communication early often need to wait for others, leaving several empty intervals of resource usage (noise). We only consider the critical duration in $\mu_{f,w}$ computation.
The definition of critical execution duration, $L(e)=[l_{c},r_{c}]$, is not trivial, and is a subinterval of $f$’s function execution duration $[l,r]$, as shown in Figure 10. This is because in practical LMT tasks, particularly within collective communication functions, numerous synchronization operations are involved. Workers that enter the collective communication kernel earlier usually perform a part of communication and then wait for others. Consequently, the throughput of network transmission are usually not continuous. The average resource utilization over the entire function execution does not accurately reflect the real communication performance. Therefore, $L(e)$ is defined as the longest interval with intense resource usage among the whole function execution duration. We apply Algorithm 1 to identify $L_{e}$ with discrete sampling of hardware resource usage.
@@
0:  Critical execution duration $[l_{c},r_{c}]$
@@
4:     $g\leftarrow$ midpoint of $[g_{\text{left}},g_{\text{right}}]$
5:     if there exists a subinterval $[l^{\prime},r^{\prime}]$ satisfying:
  1. (1)
@@
6:      Update $g_{\max}\leftarrow g$, $[l_{c},r_{c}]\leftarrow[l^{\prime},r^{\prime}]$
@@
Figure 11. The data size of a single worker’s runtime behavior patterns compared with its raw profiling data.
Figure $1$(b) shows the size of runtime behavior patterns (30 KB) is $10^{5}\times$ smaller than the raw profiling data ($\sim$3 GB). Patterns of Python functions contributes to the majority of the data volume, because PerfTracker records the full call stack of each function, which is long (sometimes the call stack of a Python function includes 1,000 letters).
4.3. Localization
PerfTracker localizes the root causes based on the runtime behavior patterns in §4.2. In LMT execution, all performance issues can be categorized into two types: (1) a common problem of all LMT workers, like hardware misconfigurations and low-efficiency code implementation, or (2) a special problem on only a part of workers, like hardware issues. Correspondingly, given function $f$ executed on worker $w$, PerfTracker defines two types of distances, (1) Distance from expectation and (2) Differential distance, to measure its abnormality.
@@ 4.3. Localization
(6) $R_{f}=[\beta_{f}^{l},\beta_{f}^{r}]\times[\mu_{f}^{l},\mu_{f}^{r}]\times[\sigma_{f}^{l},\sigma_{f}^{r}]$
$R_{f}$ is assigned based on our production experience. For example, if $f$ is a Python function, we have $R_{f}=[0,0.01]\times[0,1]\times[0,1]$, meaning that we do not expect an LMT to be bottlenecked by a Python function for ¿1% time. If $f$ is a collective communication function, we have $R_{f}=[0,0.3]\times[0,1]\times[0,1]$. As for GPU computation kernel functions, we have $R_{f}=[0,1]\times[0,1]\times[0,1]$, i.e., a GPU computation kernel function is never considered unexpected.
@@
(8) $\hat{P_{f,w}}=(\frac{\beta_{f,w}}{\max_{w\in W}{\beta_{f,w}}},\frac{\mu_{f,w}}{\max_{w\in W}{\mu_{f,w}}},\frac{\sigma_{f,w}}{\max_{w\in W}{\sigma_{f,w}}})$
@@
(9) $\Delta_{f,w}=\frac{\sum_{w^{\prime}\in W_{N}}{I(\hat{P_{f,w}},\hat{P_{f,w^{\prime}}})}}{N}$
@@
(10) $I(x,y)=\begin{cases}0&\text{if }d_{\text{Manhattan}}(x,y)<\delta,\\ 1&\text{otherwise.}\end{cases}$
@@
Localizing abnormal function executions. Based on the two distances, a function $f$ executed on worker $w$ is regarded abnormal if:
(11) $\beta_{f,w}>0.01\ \land\ (D_{f,w}>0\ \lor\ \Delta_{f,w}>M_{f}+kMAD_{f})$
@@
where $M_{f}=median\{\Delta_{f,w^{\prime}}|w^{\prime}\in W\}$ is the median value of distance to peers, and $MAD_{f}=median\{|\Delta_{f,w^{\prime}}-M_{f}|\ |w^{\prime}\in W\}$ is Median Absolute Deviation (mad, 2024), a robust measure of statistical dispersion. $k$ is set to 5.
@@
$\beta_{f,w}$¿0.01 means the function execution contributes at least 1% to end-to-end performance. Otherwise, we do not need to care about this function since the end-to-end performance does not depend much on it. For most LMT tasks, there are no more than 20 functions satisfying this requirement. $D_{f,w}>0$ indicates the runtime behavior of function $f$ is unexpected based on our production experience, and $\Delta_{f,w}>M_{f}+kMAD_{f}$ indicates the value of $\Delta_{f,w}$ is significantly larger than most other workers.
@@
As shown in §4.2, the pattern of a worker is only about 30 KB, even for an LMT with 10,000 workers, the analyzer only need to consume 300 MB data to output the final result, which is highly efficient. In our real-world deployment, PerfTracker analyzer is executed on a single CPU core.
@@
Alternatives. To localize abnormal function executions, we also tried existing clustering algorithms, including DBSCAN (Ester et al., 1996), HDBSCAN (McInnes et al., 2017), Gaussian Mixture Model (Bishop and Nasrabadi, 2006) and Mean shift (Comaniciu and Meer, 2002). They do not work well in production because they have at least one of the following limitations: (1) failing to distinguish noises and outliers, which is important in our scenario, or (2) having too many hyper-parameters, reducing their general applicability in practical cases.
5. Implementation

### v4

@@ 4 Eroica Design
Eroica is an online performance troubleshooting system for large-scale model training (LMT). It embodies the insights discussed in §3 in a production service for customers who run LMT on our GPU clusters. Eroica employs a highly efficient approach to detect performance issues by online profiling on demand, introducing little overhead on LMTs. When a performance issue is detected, Eroica automatically troubleshoots it. Eroica uses online profiling to collect fine-grained information, but it does not aggregate raw profile data of every function; instead, Eroica summarizes the runtime behavior patterns of each function at every worker node and aggregates only the concise behavior patterns. The diagnosis is performed on the aggregated behavior patterns based on the differential observability principle.
@@
Figure 6 gives an overview of Eroica. Eroica has three main components: (1) detecting performance degradation of LMT to trigger online profiling simultaneously for all workers in LMT, (2) summarizing runtime behavior patterns of each function from raw profiling data, and (3) a centralized localization algorithm that pinpoints the root-cause function based on the behavior patterns. Both (1) and (2) are deployed per worker, while (3) is global.
@@
Usage. To deploy Eroica, GPU cluster providers do not need access to the user code. The user only needs to import Eroica in the LMT code (import Eroica); no other code change is needed. Behind the line of import, Eroica initializes the detector of performance degradation, registers handler functions to start profiling, and creates a daemon process for summarizing and uploading behavior patterns. Figure 7 illustrates the output of Eroica. As a function-centric approach, Eroica pinpoints which functions on which workers are executed abnormally, and describes how they behave differently from expectation or from other workers. In Figure 7, all workers encounter slow socket receive in data loading, indicating slow storage I/O. Worker 7 is slow in collective communication, indicating a network degradation on its connected link. In addition, worker 0,1,2, and 3 execute GEMM (a CUDA kernel function [gemm]) slowly because of not fully utilizing GPUs, indicating the problem of GPU throttling.
Eroica’s output is directly fed into an AI assistant to try automated fixing: for simple code bugs, the assistant may automatically patch the code (a real case in §6.3), whereas hardware faults or complex code issues typically still require human effort to fix (e.g., replacing the problematic hardware).
Figure 7: An example of Eroica output.
4.1  Profiling on Performance Degradation
@@
Profiling an LMT task online introduces overhead. To minimize the overhead while collecting profiling data timely when performance issues occur, Eroica automatically detects performance degradation and then performs profiling simultaneously for all workers, ensuring the problematic behavior is included in the profiling data.
@@
For an LMT task, the time spent on one training iteration is the key performance metric. Eroica devises a solution to monitor the training time of each iteration without accessing user code, and then triggers online profiling if the iteration time is abnormal. Note that approaches that require accessing user data (such as logs) do not work with our usage model.
Figure 8: Eroica’s performance degradation detection.
@@
Indicators of iteration time. A PyTorch training iteration always involves several dataloader.next() calls, followed by several optimizer.step() calls (the number depends on training parameters). In this sequence of events, the duration from the first dataloader.next() to the last optimizer.step() is regarded as the duration of a complete training iteration.
@@
When it is import-ed, Eroica wraps the two functions by adding time counters. Basically, Eroica replaces the two PyTorch functions with the wrapped versions. Since the two functions are Python functions (which are not compiled before LMT executions), the replacement is done at runtime.
@@
Iteration detection. During an LMT, Eroica continuously records event sequences of dataloader.next() and optimizer.step(). After detecting $M$ (=10 in practice) identical sequences starting with dataloader.next() and ending with optimizer.step(), this sequence is defined as the training iteration sequence.
@@
Performance degradation detection. Eroica continuously monitors incoming events of dataloader.next() and optimizer.step() with the identified training iteration sequence. Each time Eroica successfully matches a complete training iteration sequence, it records the duration of that iteration. Eroica considers that performance degradation has occurred in the following two situations: (1) The average duration of the recent $N$ (=50 in practice) iterations exceeds the recent shortest iteration time by more than 5%. (2) The current training iteration sequence has not yet been fully matched, but the time elapsed since the last event received is at least $5\times$ the average iteration duration (indicating the training is blocked). If Eroica fails to match a training iteration after $K$ (=200 in practice) consecutive event receptions, it goes back to the previous iteration detection phase to redetect the training iteration sequence. While this does not often happen in common LMT programs, it significantly enhances the robustness of Eroica when users implement special functionalities in their code, to make sure the algorithm always works. Figure 8 illustrates the procedure.
Global synchronized profiling. Each LMT worker connects to an Eroica daemon. When Eroica detects a slowdown, it notifies all daemons via TCP; each daemon signals its worker to invoke a pre-registered handler, ensuring profiling runs in the LMT main thread (required by some APIs, e.g., CUPTI in Torch Profiler).
Eroica ensures synchronized profiling across workers via iteration ID: rank-0 continuously report the current iteration ID; upon a profiling trigger, the rank-0 daemon computes unified start/stop iteration IDs for profiling (with the start set a few steps ahead to ensure no worker would miss it), and other daemons periodically poll these IDs and start/stop profiling accordingly.
@@
A profiling session lasts for 20 seconds by default (configurable). We use Torch Profiler to collect execution events (Python/CPU ops, memory ops, and CUDA kernels), and nsys to sample hardware metrics at 10 kHz (GPU, DRAM, NVLink, PCIe, and network). Each daemon aggregates and collects the profiling data.
4.2  Summarizing Behavior Patterns
@@
Eroica summarizes behavior patterns of each function at the worker level. Note that Eroica only concerns functions on the critical path of LMT (Figure 9)—only these functions contribute to the end-to-end performance, despite the huge number of function executions in an LMT.
Figure 9: The critical path of LMT performance.
@@
Eroica’s solution is (1) finding the function execution events on the critical path, including GPU computation kernels, collective communication functions, memory operations, Python functions, and all other functions executed in LMT; and (2) clustering all execution events of each function (for Python functions, the entire call stack must be identical to be considered the same function), then defining several patterns to summarize the behavior of each function.
@@
Critical path. As shown in Figure 9, LMT assigns different types of function with priorities (higher is more critical): GPU compute kernels >> memory operations (e.g., malloc, memcpy) >> collective communication kernels (e.g., NCCL AllReduce) >> Python functions. A function’s execution (or a subinterval of it) is on the LMT’s critical path iff no higher-priority function is executing during that time. For Python functions, we further require that they run in the training thread (e.g., not spawned by _bootstrap) and have no executing child calls.
The rationale for such a definition is that a well-optimized LMT should try to keep GPUs busy, so in general we should focus on GPU kernel executions and function executions during GPU idle time. Therefore, all function executions are prioritized based on their relevance to GPU computation. Only when a function unrelated to GPU computation fails to overlap with GPU computation (indicating it blocks GPU computation), it should be regarded as a possible problematic function. Reducing its execution time can improve end-to-end training performance. Otherwise, it remains irrelevant to performance since it does not block GPU computation.
Note that one could craft a counterexample against Eroica, for example, by fully overlapping communication with GPU computation, making communication stalls invisible on the critical path. However, in production LMT, computation and various communications have strict dependencies. While some communication can overlap with computation, we have never seen real workloads where all communication is fully overlapped; thus it still partially appears on the critical path and can be detected by Eroica.
@@
[…]
@@
$\mu_{f,w}$ represents the average hardware resource utilization during all executions of function $f$ on worker $w$. In this context, “resource” refers to hardware resources that determine performance of function $f$. For example, GPU computation kernels correspond to GPU SM frequency, Python functions to CPU utilization, intra-worker collective communication to NVLink utilization, and inter-worker collective communication corresponds to the bandwidth utilization between GPU and NIC, due to the significantly lower NIC bandwidth compared to intra-worker NVLink bandwidth.
$\mu_{f,w}=\frac{\sum_{e\in E_{f,w}}{|L(e)|avg_{U}(L(e))}}{\sum_{e\in E_{f}}{|L(e)|}}$ (4)
@@
Figure 10: Critical duration of a collective communication function. A worker enters the communication early often need to wait for others, leaving several empty intervals of resource usage (noise). We only consider the critical duration in $\mu_{f,w}$.
The definition of critical execution duration, $L(e)=[l_{c},r_{c}]$, is not trivial, and is a subinterval of $f$’s function execution duration $[l,r]$, as shown in Figure 10. This is because in practical LMT tasks, particularly within collective communication functions, numerous synchronization operations are involved. Workers that enter the collective communication kernel earlier usually perform a part of communication and then wait for others. Consequently, the throughput of network transmission are not continuous. The average resource utilization over the entire function execution does not accurately reflect the real communication performance. Thus, $L(e)$ is defined as the longest interval with intense resource usage among the whole function execution duration. We apply Algorithm 1 to identify $L_{e}$ with discrete sampling of hardware resource usage.
@@
0:  Critical execution duration $l_{c},r_{c}]$
@@
4:   $g\leftarrow$ midpoint of $g_{\text{left}},g_{\text{right}}]$
5:   if there exists a subinterval $l^{\prime},r^{\prime}]$ satisfying:
  1. 1.
@@
6:     Update $g_{\max}\leftarrow g$, $l_{c},r_{c}]\leftarrow[l^{\prime},r^{\prime}]$
@@
Note that all three dimensions of $P_{f,w}$ are independent of absolute timestamps. For example, to compute $\beta_{f,w}$, we only need the time difference between the start and end of function execution. This design is the key to performing inter-host pattern comparison without the need for clock synchronization.
@@
Figure 11: Data size of a single worker’s runtime behavior patterns compared with its raw profiling data.
Figure 1111(b) shows the size of runtime behavior patterns (30 KB) is $10^{5}\times$ smaller than the raw profiling data ($\sim$3 GB). Patterns of Python functions contributes to the majority of the data volume, because Eroica records the full call stack of each function, which is long (sometimes the call stack of a Python function includes 1,000 letters).
4.3  Localization
Eroica localizes the abnormal function behavior based on the runtime behavior patterns in §4.2. In most cases, the abnormal function behavior can directly pinpoint a single plausible root cause. For more complex cases, our engineers make the final call, optionally with the help of an AI assistant.
In LMT execution, all performance issues can be categorized into two types: (1) a common problem of all LMT workers, like hardware misconfigurations and low-efficiency code implementation, or (2) a special problem on only a part of workers, like hardware issues. Correspondingly, given function $f$ executed on worker $w$, Eroica defines two types of distances, (1) Distance from expectation and (2) Differential distance, to measure its abnormality.
@@ 4.3  Localization
$R_{f}=[\beta_{f}^{l},\beta_{f}^{r}]\times[\mu_{f}^{l},\mu_{f}^{r}]\times[\sigma_{f}^{l},\sigma_{f}^{r}]$ (6)
$R_{f}$ is assigned based on our production experience. For example, if $f$ is a Python function, we set $R_{f}=[0,0.01]\times[0,1]\times[0,1]$, since customers typically regard performance fluctuations within 1% as measurement noise; deviations beyond this threshold are viewed as substantial regressions and thus trigger diagnostic requests. If $f$ is a collective communication function, we set $R_{f}=[0,0.3]\times[0,1]\times[0,1]$. For GPU computation kernel functions, we set $R_{f}=[0,1]\times[0,1]\times[0,1]$, i.e., we expect a LMT to spend most of its time on GPU computation to fully utilize the GPUs.
@@
$\hat{P_{f,w}}=(\frac{\beta_{f,w}}{\max_{w\in W}{\beta_{f,w}}},\frac{\mu_{f,w}}{\max_{w\in W}{\mu_{f,w}}},\frac{\sigma_{f,w}}{\max_{w\in W}{\sigma_{f,w}}})$ (8)
@@
$\Delta_{f,w}=\frac{\sum_{w^{\prime}\in W_{N}}{I(\hat{P_{f,w}},\hat{P_{f,w^{\prime}}})}}{N}$ (9)
@@
$I(x,y)=\begin{cases}0&\text{if }d_{\text{Manhattan}}(x,y)<\delta,\\ 1&\text{otherwise.}\end{cases}$ (10)
@@
Localizing abnormal function executions. With the two distances, a function $f$ running on worker $w$ is abnormal if:
$\beta_{f,w}>0.01\ \land\ (D_{f,w}>0\ \lor\ \Delta_{f,w}>M_{f}+kMAD_{f})$ (11)
@@
where $M_{f}=median\{\Delta_{f,w^{\prime}}|w^{\prime}\in W\}$ is the median value of distance to peers, and $MAD_{f}=median\{|\Delta_{f,w^{\prime}}-M_{f}|\ |w^{\prime}\in W\}$ is Median Absolute Deviation [mad], a robust measure of statistical dispersion. $k$ is set to 5.
@@
$\beta_{f,w}$>0.01 means the function execution contributes at least 1% to end-to-end performance. For most LMT tasks, no more than 20 functions satisfy this requirement. $D_{f,w}>0$ indicates the runtime behavior of function $f$ is unexpected based on our production experience, and $\Delta_{f,w}>M_{f}+kMAD_{f}$ indicates the value of $\Delta_{f,w}$ is significantly larger than most other workers.
@@
As shown in §4.2, the pattern of a worker is only $\sim$30 KB, even for an LMT with 10,000 workers, the localization algorithm only consumes 300 MB data to output the final result, which is highly efficient to be executed on a single CPU core.
@@
Alternatives. To localize abnormal function executions, we also tried existing clustering algorithms, including DBSCAN [ester1996density], HDBSCAN [mcinnes2017hdbscan], Gaussian Mixture Model [bishop2006pattern] and Mean shift [comaniciu2002mean]. They do not work well in production because they have at least one of the following limitations: (1) failing to distinguish noises and outliers, which is important in our scenario, or (2) having too many hyper-parameters, reducing their general applicability in practical cases.
5 Implementation